# Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation

### v1

@@ 2.2 Results
Effect of Negation Argument. To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options. This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C." (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C)." As shown in Table 2, all the MLLMs show higher accuracy for the second strategy after negation. The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments. Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments. Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods. Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
3 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA li2024llavanext-strong , Qwen2-VL Qwen2VL-7B and Deepseek liu2024deepseek , have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development . In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert . To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding . This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation. These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts. Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can . Very recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments. Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
In multimodal learning, negation has received comparatively less attention than in NLP. Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision . For instance, very recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance. In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
@@ 4 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in their alignment and reasoning mechanisms. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.

### v2

@@ 4.2 Discussion: Effect of Negation Argument
To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options. This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C." (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C)." As shown in Table 3, all the MLLMs show higher accuracy for the second strategy after negation. The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments. Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments. Furthermore, as shown in Figure 8, the choice of negation arguments significantly affects MLLMs’ performance. For instance, when the negation argument includes “ten" or “eight," only two models maintain the correct response of “four," whereas using “one" as the negation argument allows six models to retain the correct answer. This highlights the varying sensitivity of MLLMs to different negation formulations. Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods. Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
@@ 5 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in their alignment and reasoning mechanisms. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial negation conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.